Part A How long does it take the Sun to melt a block of ice at...

Part A

How long does it take the Sun to melt a block of ice at 0∘C with a flat horizontal area 1.0 m2 and thickness 2.0 cm ? Assume that the Sun's rays make an angle of 44 ∘ with the vertical and that the emissivity of ice is 0.050.

Express your answer to two significant figures and include the appropriate units.

2.12min , or 127.7s, or 130s incorrect
Answer #1

Q to melt the this ice = m (ice)* Lf
= density of ice * volume of ice * latent heat of ice
= density of ice* flat area of ice * thickness of ice * latent heat of ice
= 917 g/m^3 * 1 m^2 * (2 * 10^-2 m) * 334 J/g
=6125.56 J
=6.126 KJ


now the sun rays makes an angle of 44 degree with the vertical so the angle made with horizontal =46 degree
solar maximum flux = 1.362kW/m^2 (when perpendicular to surface)
I = 1.362 * cos(46) = 0.946 KW/m^2
net heat absorb by ice = emissivity of ice * intensity of sun * area of ice *t
6.126 =0.05 * 0.946 * 1 *t
=130 seconds